Please help me solve this equation: Point R is on line segment PQ such that PR=35 and PR:RQ = 5:6. Find RQ.
Answer by amoresroy(361)

PR/RQ = 5/6
RQ = 6/5(PR)
RQ = 6/5(35)
RQ = 42